OpenStudy (anonymous):

Please help me!? What is the solution of x^2+6x=-5?

OpenStudy (anonymous):

First, add 5 to both sides of the equation. Now it becomes x^2+6x+5=0 Now factor this equation. (x+2)(x+3) = 0 The answer is therefore x=-2,-3

OpenStudy (anonymous):

Well My problem says x^2+6x=-5 and its multiple choice. 1) -6 2) -1 3) 1 4)6

OpenStudy (anonymous):

Oops! I just woke up, I totally factored that wrong. It is (x+1)(x+5), so x = -1,-5 Pick 2) Check this out if you want to learn how to factor. http://www.mathsisfun.com/algebra/quadratic-equation.html

OpenStudy (anonymous):

Okay thank you! Could you see if i got my other 3 answers correct?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1.) what is the value of n so that the expression is x^2+11x+n is a perfect square trinomial? 11 25 30.25(my answer) 36 3.)Which of the following is a solution of x^2+14x+112=0? If necessary, round to the nearest hundredth. x= -0.24 x= -4.24 4.24 no solution(my answer) 4)A box shaped like rectanglular prism has a height of 17 in and a volume of 2,700 in^3 the length is 4 inches greater than twice the width. What is the width of the box? -10 in 8 in (my answer) 20 in 40 in

OpenStudy (anonymous):

Nevermind i got 100% :) Number 2 was -1

OpenStudy (anonymous):

okay cool.

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (samslays):

Yep 100% I just took this test